# PROPS: Progressively Private Self-alignment of Large Language Models

### v1

@@ A.6 Additional Experimental Results
PROPS vs DP-SGD and PROPS vs RR: We first present supplementary experimental results and their corresponding standard derivations. Specifically, we report means and standard deviations of the win-tie rates of PROPS vs. RR and PROPS vs. DP-SGD in Tables 6 and 7 respectively. Unlike DP-SGD, which employs gradient perturbation, PROPS utilizes an input perturbation mechanism that maintains the post-processing property of differential privacy. This inherent flexibility enables extensive hyper-parameter tuning, including training epochs, without compromising privacy guarantees. Consequently, despite fluctuations in the win-tie-loss ratio as determined by the GPT-4 model, we still observe consistent trends. Notably, in the high privacy regime (i.e., a low privacy budget), PROPS exhibits a higher win rate. Conversely, in scenarios with a more relaxed privacy constraint, DP-SGD demonstrates better performance. Another noteworthy discrepancy arises from the interplay between model size and model family as larger models benefit from PROPS more. This discrepancy may be attributed to the different learning capacities of the model under varying privacy requirements. Specifically, when the preference dataset is subjected to consistent noise injection, the larger models in the initial training stages exhibit better learning, thereby positively influencing subsequent model iterations.
@@ A.7 Win-rate Evaluation & Additional Prompt-Response Pair Examples as a Function of Privacy Budget
Win-rate Evaluation: We used GPT4 to obtain win-tie rate measurements by comparing responses generated by PROPS against conventional privacy-based alignment methods (DP-SGD and RR). Figure 6 shows an example of how these measurements were obtained. We instruct GPT4 to act as an evaluator/judge (to assign the Helpfulness and Harmlessness scores) by providing it the following prompt similar to Chakraborty et al. (2024), “You are a helpful, harmless, and precise assistant for checking the quality of the answer. We would like to request your feedback on the performance of two AI assistants in response to the user question. Please rate the harmlessness, helpfulness, and level of detail of their responses. Your evaluation should consider factors such as the helpfulness, harmlessness, relevance, accuracy, depth, creativity, and level of detail of the response. Bear in mind that the response might be truncated at the end due to length constraints. Each assistant receives an overall score on a scale of 1 to 10 (or 1 to 5), where a higher score indicates better overall performance. Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.” For the same prompt, “Do violent video games cause people to become more violent in real life”, as shown in Figure 6, we provide GPT4 responses generated by GPT2-Large trained under PROPS and DP-SGD. As the figure shows, GPT4 assigns a higher score to the PROPS response, ultimately giving it a win for this prompt.
@@
Figure 9 shows two prompt-response examples generated by PROPS, DP-SGD, and RR. Similar trends are observed between PROPS and DP-SGD as in the previous examples, where DP-SGD generates similar responses across privacy budgets while PROPS improves in response quality when the privacy budget is increased. RR also seems to exhibit similar behaviors where for the prompt, “Can humans catch warts from frogs or toads?”, it generates a more focused response at $\epsilon=0.1$ rather than $\epsilon=1$. However, looking at $\epsilon=0.1$, the figure indicates that PROPS is able to generate an answer that better addresses the concern of the user compared to vanilla RR.

### v2

@@ A.6 Additional Experimental Results
PROPS vs DP-SGD and PROPS vs RR: We first present supplementary experimental results and their corresponding standard derivations. Specifically, we report means and standard deviations of the win-tie rates of PROPS vs. RR and PROPS vs. DP-SGD in Tables 7 and 8 respectively. Unlike DP-SGD, which employs gradient perturbation, PROPS utilizes an input perturbation mechanism that maintains the post-processing property of differential privacy. This inherent flexibility enables extensive hyper-parameter tuning, including training epochs, without compromising privacy guarantees. Consequently, despite fluctuations in the win-tie-loss ratio as determined by the GPT-4 model, we still observe consistent trends. Notably, in the high privacy regime (i.e., a low privacy budget), PROPS exhibits a higher win rate. Conversely, in scenarios with a more relaxed privacy constraint, DP-SGD demonstrates better performance. Another noteworthy discrepancy arises from the interplay between model size and model family as larger models benefit from PROPS more. This discrepancy may be attributed to the different learning capacities of the model under varying privacy requirements. Specifically, when the preference dataset is subjected to consistent noise injection, the larger models in the initial training stages exhibit better learning, thereby positively influencing subsequent model iterations.
@@ A.7 Win-rate Evaluation & Additional Prompt-Response Pair Examples as a Function of Privacy Budget
Win-rate Evaluation: We used GPT4 to obtain win-tie rate measurements by comparing responses generated by PROPS against conventional privacy-based alignment methods (DP-SGD and RR). Figure 6 shows an example of how these measurements were obtained. We instruct GPT4 to act as an evaluator/judge (to assign the Helpfulness and Harmlessness scores) by providing it the following prompt similar to chakraborty2024transferqstarprincipled, “You are a helpful, harmless, and precise assistant for checking the quality of the answer. We would like to request your feedback on the performance of two AI assistants in response to the user question. Please rate the harmlessness, helpfulness, and level of detail of their responses. Your evaluation should consider factors such as the helpfulness, harmlessness, relevance, accuracy, depth, creativity, and level of detail of the response. Bear in mind that the response might be truncated at the end due to length constraints. Each assistant receives an overall score on a scale of 1 to 10 (or 1 to 5), where a higher score indicates better overall performance. Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment." For the same prompt, “Do violent video games cause people to become more violent in real life", as shown in Figure 6, we provide GPT4 responses generated by GPT2-Large trained under PROPS and DP-SGD. As the figure shows, GPT4 assigns a higher score to the PROPS response, ultimately giving it a win for this prompt.
@@
Figure 9 shows two prompt-response examples generated by PROPS, DP-SGD, and RR. Similar trends are observed between PROPS and DP-SGD as in the previous examples, where DP-SGD generates similar responses across privacy budgets while PROPS improves in response quality when the privacy budget is increased. RR also seems to exhibit similar behaviors where for the prompt, “Can humans catch warts from frogs or toads?", it generates a more focused response at $\epsilon=0.1$ rather than $\epsilon=1$. However, looking at $\epsilon=0.1$, the figure indicates that PROPS is able to generate an answer that better addresses the concern of the user compared to vanilla RR.